OpenStudy (anonymous):

the measure of the supplement of an angle is 20 degrees more than three times the original angle . Find the measures of the angles.

OpenStudy (nurali):

Let x be the measure of original angle, then the measure of the supplement is 3x + 20. We can set up an equation using the fact that the sum of an angle and its supplement is 180. So x + (3x + 20) = 180 Solving the equation for x, we have x + 3x + 20 = 180 4x = 160 x = 40 So the measures of the angles are 40 and (180 - 40 =) 140 degrees.

OpenStudy (campbell_st):

one angle = x the other is 3x + 20 supplementary means they add to 180 so x + 3x + 20 = 180 solve for x

OpenStudy (anonymous):

thanks guys

OpenStudy (nurali):

My Pleasure.